# Strategic Planning and Rationalizing on Trees Make LLMs Better Debaters

### v1

@@ Appendix C Human Evaluation Details
Motion List We collect 52 debate motions from different sources, including the recent hot topics in the debate website OpentoDebate [^5] https://opentodebate.org/ and the opinion section in the New York Times, the motions used in previous debate systems (Slonim et al., 2021; Zhang et al., 2024), and the persuasiveness dataset released by Anthropic (Durmus et al., 2024). It covers various domains such as Economics, Finance, Health, Science, Culture, and etc. We then asked two human expert debaters to annotate how polarized the motions are based on a 1-5 scale. Finally, we keep 13 motions that are less polarized. We list motions and their sources in Table 4. Note that these topics are only used as a debate motion. We do not use the background materials or debate transcripts in their sources.
@@ 9 Artists should be free to borrow from cultures other than their own Culture OpentoDebate
Recruitment We recruit audience members from the online research platform Prolific  [^6] https://www.prolific.com/. The evaluation takes around 30 minutes per stage. The compensation is \$10 for each case after they complete the evaluation. All participants are located in the US when they conduct the evaluation. Compensation is higher than the minimum wage of \$7.25 per hour in the United States. The participants should be older than 18 years old, be fluent in English, and have achieved a high school degree. We have two attention check questions during the evaluation to control the quality. We remove the data points that fail the check (about 10%). The screenshot of our evaluation platform is shown in Figure 5 and 6.
@@ Appendix D Human Debate Flow Tree Corpus
We create the human debate flow tree corpus from two debate datasets: PanelBench (Liang et al., 2024). PanelBench is built from the online debate platforms DebateArt and world-class competitive debate competitions, BP-Competition. The number of debate is 122. We follow the MIT license of PanelBench. We first reorganize the original debate transcript into the opening, rebuttal, and closing statements. Then, for each statement, we extract the debate action tuples (action, claim, argument, target claim) and build the debate flow tree based on Alg. 2 for each debate. These human debate flow trees will then be retrieved during the debate to provide feedback for TreeDebater. There is no overlap between the motions used in our experiments with the human debate flow tree corpus.
To endow TreeDebater with sophisticated, human-like debating capabilities, we incorporate the human debate tree into the pipeline. As highlighted in Section 1, human experts intuitively employ tree-like structures to organize potential arguments and to meticulously track the evolving flow of a debate. Our human debate feedback mechanism operationalizes this human-centric approach. The core of this mechanism lies in comparing TreeDebater’s current debate flow tree with a curated corpus of human debate flow trees. By employing semantic search based on their embeddings (Reimers & Gurevych, 2019), the system identifies human debates that exhibit the most similar argumentative structures to the ongoing debate.
@@ Appendix F Additional Experimental Results
The standard deviation of Table 1 and 2 are shown in Table 6 and 7.
@@ Appendix G Main Prompts
## Workflow
1. Create a opening plan based on the current debate flow trees. It should include the definition, judging criteria, and the battlefields to discuss.
2. Select among the definition, judging criteria, and battlefields and allocate your word budget based on the importance of each point. You should allocate more words to the more important points and can ignore the less important points.
3. Follow the opening plan to generate conversational arguments.
4. Deliver a {n_words}-word opening statement.
@@
Your Debate Tree:
* Level-1 Nodes: Your main claims and arguments
* Level-2 Nodes: Opponent’s attacks on your claims and arguments
* Level-3 Nodes: Your rebuttal on the attacks
Opponent’s Debate Tree:
* Level-1 Nodes: Opponent’s main claims and arguments
* Level-2 Nodes: Your attacks on their claims and arguments
* Level-3 Nodes: Opponent’s rebuttal on your attacks
@@
Your Tree: {tree}
Opponent’s Tree: {oppo_tree}
Your Main Claims: {claims}
Definition: {definition}
@@
Table 9: Opening Stage Generation Prompt.
@@
## Workflow:
1. Analyze the debate flow trees to select the battlefields you will discuss from the given list of battlefields.
2. Allocate your word budget among the selected battlefields based on their importance and create the rebuttal plan. The plan should include the overview of battlefields you will discuss and the word budget for each battlefield. You should allocate more words to the more important points and can ignore the less important points.
3. Follow the rebuttal plan to generate conversational arguments.
4. Deliver a rebuttal statement with {n_words} words in total. Present only the final text.
@@
Your Debate Tree:
* Level-1 Nodes: Your main claims and arguments
* Level-2 Nodes: Opponent’s attacks on your claims and arguments
* Level-3 Nodes: Your rebuttal on the attacks
Opponent’s Debate Tree:
* Level-1 Nodes: Opponent’s main claims and arguments
* Level-2 Nodes: Your attacks on their claims and arguments
* Level-3 Nodes: Opponent’s rebuttal on your attacks
@@
Your Tree: {tree}
Opponent’s Tree: {oppo_tree}
@@
## Workflow
1. Analyze the debate flow trees to strategically allocate your word budget based on argument importance and status. Structure your response according to the above points.
2. Generate conversational yet language-suitable arguments following the allocated word budget.
@@
Your Tree:
* Level-1: Your main claims and arguments
* Level-2: Opponent’s rebuttals on your main claims and arguments
* Level-3: Your defenses on the rebuttals
Opponent’s Tree:
* Level-1: Opponent’s main claims and arguments
* Level-2: Your rebuttals on the opponent’s main claims and arguments
* Level-3: Opponent’s defenses on your rebuttals
@@
Your Tree: {tree}
Opponent’s Tree: {oppo_tree}
@@ 3. Issue: The judging criterion justification can be strengthened
Table 12: Example of Human Debated Based Feedback.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main claims are also the key points of the main paper, including the method and experiment sections.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, see the last section in our paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: There is no theoretical results in our paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide most of the relevant details for implementation in our paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the code and dataset in the supplemental material
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We train reward models in our paper and include the training details in the appendix
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide the standard deviation of our main results in the appendix
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We can run our experiments with one A6000 GPU and tens of CPU cores in the cluster. The running time does not depend much on GPU resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We try our best to adhere to the NuerIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, see the last section in our paper.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release data or models that might have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original paper of the datasets and provide the version and license of each dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the anonymized code and its documents in the supplemental materials.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We got our IRB approval and include the full text of instructions in the appendix.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: We have IRB approvals and are adhering to the study rules in our submitted files.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this project does not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ Appendix C Human Evaluation Details
Motion List We collect 52 debate motions from different sources, including the recent hot topics in the debate website OpentoDebate [^7] https://opentodebate.org/ and the opinion section in the New York Times, the motions used in previous debate systems [undefad, undefal], and the persuasiveness dataset released by Anthropic [undefg]. It covers various domains such as Economics, Finance, Health, Science, Culture, and etc. We then asked two human expert debaters to annotate how polarized the motions are based on a 1-5 scale. Finally, we keep 13 motions that are less polarized. We list motions and their sources in Table 5. Note that these topics are only used as a debate motion. We do not use the background materials or debate transcripts in their sources.
@@ 9 Artists should be free to borrow from cultures other than their own Culture OpentoDebate
Recruitment We recruit audience members from the online research platform Prolific  [^8] https://www.prolific.com/. The evaluation takes around 30 minutes per stage. The compensation is \$10 for each case after they complete the evaluation, which is higher than the minimum wage of \$7.25 per hour in the United States. The participants should be older than 18 years old, be fluent in English, and have achieved a high school degree. We have two attention check questions during the evaluation to control the quality. Two screening questions are : (i) a multiple-choice question to choose the main claim proposed by a specific side, with multiple confusing options that are difficult to distinguish; (ii) a free-form QA question to summarize the key idea of one side. We filter out responses that fail the screening questions (about 10%). One example is shown below. The screenshot of our evaluation platform is shown in Figure 5 and 6. The demographic survey results of our participants are shown in Figure 7.
Two screening questions Q1: Which claim is proposed by For side as its first main claim during the opening statement? If multiple options apply, please choose the best one.
(A) The debt ceiling creates unnecessary political crises that harm the economy.
(B) Effective Altruism’s metrics-driven approach overlooks crucial local contexts that determine real impact
(C) The debt ceiling undermines the full faith and credit of the United States.
(D) The debt ceiling encourages bipartisan negotiation on fiscal policy.
(E) None of the above
Q2: Which is the main battlefield / conflict / question mentioned by For side in its closing statement? *
@@ Appendix D Human Debate Flow Tree Corpus
We create the human Debate Flow Tree corpus from the human debate dataset PanelBench [undefr]. PanelBench is built from the online debate platforms DebateArt and world-class competitive debate competitions, BP-Competition. The number of debate is 122. We follow the MIT license of PanelBench. We first reorganize the original debate transcript into the opening, rebuttal, and closing statements. Then, for each statement, we extract the debate action tuples (action, claim, argument, target claim) via the backbone LLM and build the Debate Flow Tree based on Alg. 2 for each debate. These human Debate Flow Trees will then be retrieved during the debate to provide feedback for TreeDebater. There is no overlap between the motions used in our experiments with the human Debate Flow Tree corpus.
To endow TreeDebater with sophisticated, human-like debating capabilities, we incorporate the human debate tree into the pipeline. As highlighted in Section 1, human experts intuitively employ tree-like structures to organize potential arguments and to meticulously track the evolving flow of a debate. Our human debate feedback mechanism operationalizes this human-centric approach. Specifically, during the debate, we reformulate the current Debate Flow Tree into a tree-like string. Then we employ semantic search based on the embeddings of the debate flow tree string. One of the state-of-the-art embedding models Gemini-text-embedding-4 is used to get the embedding of the tree-like string of the current debate flow tree and of trees in human debate flow tree corpus. The top 1 human debate flow tree is retrieved with a similar threshold of 0.8. We add the retrieval to the instruction for simulated audience to let it aware of how human debate on similar topics.
@@ Appendix F Additional Experimental Results
The standard deviation of Table 1 is shown in Table 7.
We conducted 10 additional evaluations for each stage, providing a more robust statistical foundation. As shown below, our method consistently outperforms baselines across 5 head-to-head stages in stage-level win rate, with comparable performance in the remaining 1 stage. Among the 5 stages, 3 of them (Opening Pro, Rebuttal Pro, and Closing Con) showed significant improvements (95% CI, p<0.05). Critically, our method shows significant overall performance (95% CI, p<0.05). Notably, in debate settings, it is almost unlikely for one debater to win all stages without ties, even for human debate experts. Considering the overall performance, the consistently better win rate with significant overall gains represents strong empirical evidence of our method’s effectiveness. We will add this statistical reporting, along with the standard deviations, to our revised manuscript.
Table 9: Statistical Significance. Stages with * indicate statistical significance, which are demonstrated by the positive confidence interval and p-value < 0.05.
Our Win Rate Confidence Interval (CI 95%) P-value
Opening Pro* 88.90% (0.225, 1.047) 0.002
Opening Con 57.10% / >0.05
Rebuttal Pro* 70.00% (0.054, 0.612) 0.019
Rebuttal Con 57.10% / >0.05
Closing Pro 50.00% / >0.05
Closing Con* 85.70% (0.104, 0.646) 0.006
Overall* 71.73% (0.138, 0.459) 0.0002
@@ Appendix G Main Prompts
Main Claim Generation Prompt
## Task: Generate Strategic Counter-Arguments
You are participating in a formal debate on the motion: {motion}
Your position: {act} the motion
## Your Objective
Generate num persuasive counter-arguments that:
## Context
Previous debate exchanges:
{history}
Table 10: Main Claim Generation Prompt.
@@
## Rules
- Ensure your language is fluent and natural, indistinguishable from human writing. Make sure your debate script is complete, including definitions, judging criteria, and arguments.
- When citing data and academic research, provide sources within the context and avoid using information not present in the provided materials. Ensure your arguments are supported by data and academic evidence.
- When citing data, **using specific figures** instead of just descriptive language will make your argument more persuasive.
- When citing data and academic research, **don’t just** list the information, **but also explain** how it supports your point.
- When choosing evidence, numeric data is preferred over qualitative descriptions.
- List your references in a standard format in a Reference section, making sure each source has a clear number and url information.
- Renumber the original citations to [1],[2], … and etc. List full references in **Reference** section. This section should start with **Reference**. Use Chicago style for the reference list and DO NOT include the web link or url information.
@@
Your Tree: {tree} Opponent’s Tree: {oppo_tree} Your Main Claims: {claims} Definition: {definition}
@@
Table 12: Opening Stage Generation Prompt. Instructions marked with blue are borrowed from Agent4Debate.
@@
## Knowledge
### Structure of a Rebuttal
A complete rebuttal should consist of multiple points, with each point containing four parts:
- **Lead-in:** Introduce the opponent’s argument, evidence, reasoning, etc., that you will be refuting.
- **Explanation:** Briefly explain the opponent’s argument to ensure clarity.
- **Rebuttal:** This is the core of your point. Utilize rebuttal techniques to directly challenge the opponent’s claim.
- **Impact:** Concisely summarize the impact of your rebuttal and how it benefits your side.
Note: Typically, the lead-in and explanation are combined into one sentence. The rebuttal is the most crucial part, and the impact summarizes its effect.
### Rebuttal Techniques
- **Pointing out logical fallacies:** Identify errors in the opponent’s reasoning, such as reversing cause and effect, equivocation (shifting the meaning of a key term), straw man arguments, circular reasoning, or tautology (repeating the same idea in different words).
- **Pointing out factual errors:** Highlight inaccuracies or weaknesses in the opponent’s evidence, such as insufficient data, incorrect facts, or biased sources.
- **Pointing out error logic:** Identify flawed logic underlying opponent’s framework.
- **Leveling the playing field:** This technique aims to neutralize the opponent’s advantage or minimize the perceived harm of your side’s position by demonstrating that both sides share the same issue or benefit. - **Acknowledging and countering:** Start by acknowledging a valid point made by your opponent before explaining why your position still offers a better solution.
@@
Your Tree: {tree} Opponent’s Tree: {oppo_tree}
@@
## Primary Objectives of a Closing Statement
- Convince the judges that your team won more battlegrounds.
- Demonstrate your team’s strengths and the opponent’s weaknesses within each battleground based on the clash outcomes.
## Rules - Ensure your writing is fluent, natural, and indistinguishable from human writing.
- Avoid empty appeals to values. Remember, value appeals should connect back to the topic and your stance.
- When citing data or theories, provide sources. Do not introduce new information or data in the closing statement.
- Avoid repeating arguments from previous speeches. Instead of mechanically listing points, focus on deepening your arguments and ensuring logical coherence.
- This is a closing statement, not the time for new arguments. Prioritize depth over breadth.
- Base your arguments on the identified battlegrounds and clashes.
@@
Your Tree: {tree} Opponent’s Tree: {oppo_tree}
@@ 3. Issue: The judging criterion justification can be strengthened
Table 15: Example of Human Debated-Based Feedback.
Simulated Audience Feedback Prompt
## Your Task
You are a panel of debate audience members to provide comprehensive feedback on how the statement impacts and persuades a general audience.
### Audience Panel Composition
- General public with varied educational backgrounds
- Students and educators from different fields
- Professionals interested in policy and social issues
### Evaluation Dimensions
1. **Core Message Clarity**
2. **Engagement Impact**
3. **Evidence Presentation**
4. **Persuasive Elements**
### Guidelines
- Evaluate all dimensions thoroughly
- Identify the most significant barriers to audience understanding in the {stage} statement
- Consider which issues could be addressed with minimal revisions on the {stage} statement
- Focus on high-impact, low-disruption improvements
## Retrieval Information
Here are debate flow trees and action allocations from human debates.
Use the structure and allocation strategy to provide better feedback.
{retrieval debate flow tree}
### Input Information
Debate Topic: {motion}
History of the debate: {history}
Current {side}’s {stage} Statement to be evaluated: {statement}
### Output Format
[Comprehensive Analysis]
Core Message Clarity:
Engagement Impact:
Evidence Presentation:
Persuasive Elements:
[Critical Issues and Minimal Revision Suggestions]
Issue:
Impact on Audience:
Minimal Revision Suggestion:
Table 16: Simulated audience feedback prompt.